# Influence of finite-temperature effects on CMB power spectrum

### v1

@@ 3 Modified CLASS and machine learning techniques
Weinberg’s approach is insightful and valuable for understanding the physics of key events during the intermediate period between the generation of the CMB and its observation today. However, as noted in his book, its primary aim is not the precise calculation of the power spectrum curve. For such purposes, computational tools like CAMB or CLASS are typically employed. In this section, we utilize a finite-temperature modified version of CLASS and perform two types of analyses. The first analysis involves a brute-force scan, as described below. The second analysis complements the first by incorporating statistical and machine-learning techniques.
@@
$H=7.204\times 10^{-19}\,T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}\,T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}+\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}{\color[rgb]{0,0,1}+\frac{2.725}{T}\Omega_{\Lambda_{3}}h^{2}}}.$
Once the values of the cosmological parameters are fixed, the power spectra are computed using the modified CLASS. A large number of parameter sets can be scanned by running the CLASS Python wrapper, CLASSy. Each resulting spectrum is then fed into a code that calculates the Euclidean distance between the generated curve and the Planck data [3]. By these comprehensive scans and distance calculations we obtain tens of millions of data points. The details of the brute-force scan will vary slightly each time because the data collection procedure involves random sampling. It will be nice to check the qualitative features of the results by an independent method. For this purpose we employ machine learning techniques and evaluate the predictive performance of the models in estimating the distance.
We consider two cosmological models: the first is the original $\Lambda$CDM model, which has 7 parameters: $(\Omega_{K},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{\text{reio}})$. The second model incorporates finite-T QG effects and has 8 parameters: $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{\text{reio}})$. For each model, we use 5000 datasets, selected from a much larger data sample as explained below. We then apply a regression-based machine learning technique to estimate the minimum distance between the generated spectra and the Planck 2018 data.
The objective is to determine which model best explains the variance in the target variable, the distance $d$, while balancing model complexity and performance. Model selection is based on maximizing $R^{2}$ and minimizing the MSE in both training and testing datasets. Using quartic regression, which proves to be the best-performing model, we compare the accuracies and generalization capabilities of both models. The comparison shows that including the additional parameters from finite-temperature QG effects modifies the values of the cosmological parameters, resulting in an improved distance fit to the Planck 2018 data.
@@ 3.1 Brute-force scan by CLASSy
In the analysis by employing CLASS, we first conduct brute-force scans to systematically explore the parameter space for cosmological models. By generating discrete power spectra over the multipole moment parameter $\ell$ for suitable ranges of the parameters, their distances from the Planck 2018 experimental curve can be calculated. This approach requires extensive scanning of the parameter space, ultimately determining the set of cosmological parameters that minimizes the distance to the Planck 2018 curve. The brute-force method demands high-performance computing resources. Using CLASS, specifically its Python wrapper CLASSy, we conduct extensive parameter scans, ultimately collecting several tens of millions of datasets. This immense volume of data is necessary to perform various checks and validate the robustness of the results. Based on these preliminary explorations, appropriate parameter ranges are carefully refined. From these final intervals, we collect 400,000 datasets for the 7-parameter case and 300,000 datasets for the 8-parameter case. The number of the datasets for the 8-parameter model was deliberately chosen less than that of the 7-parameter model. This was to bring out the fact that the former performs better in spite of this disadvantage. See the related comments below. These tasks test the limits of our high-end personal computers, which, while capable, are pushed to their full capacity. (Despite leveraging a supercomputer for part of the data collection, the additional computational power offered minimal speed advantages, emphasizing the intense demands of such exhaustive analyses. See the conclusion for more.)
The impact of inclusion of $\Omega_{\Lambda_{2}}$ is reflected on the actual minimum values of the distance and the corresponding parameters set. For the 7-parameter model, the minimum distance of the 400,000 datasets is
@@
$\Omega_{\Lambda_{2}}$ $\Omega_{\Lambda_{3}}$ $h$ $\omega_{b}$ $\omega_{cdm}$ $A_{s}$ $n_{s}$ $\tau_{reio}$
$-2.073444\times 10^{-8}$ 0.004297 0.675500 0.022394 0.119889 $2.101111\times 10^{-9}$ 0.966000 0.054444
Note that these minima are actual data values. Below we obtain the regression functions of these two datasets and obtain the minima (given in the ”Range” colums of the table in Fig. 4) predicted by the functions. Those minima will in turn be checked against the values (given in (33)) yielded by directly running CLASS with the corresponding values of the parameters as input. To provide some context for the distance values, we show two plots in Fig. 3.As seen in these plots, the curves with distances in the range of several hundreds are nearly identical to the Planck 2018 curve, appearing directly on top of it.
@@
The analysis of the 8-parameter model, based on 300,000 datasets, compared to the 400,000 datasets used for the 7-parameter model, highlights an inherent limitation in the dataset sizes and their comparability. Given that the 8-parameter model includes one additional parameter, a fair comparison would require a dataset larger than 400,000 by some appropriate factor to adequately cover the expanded parameter space. Despite this discrepancy, the results indicate that the minimum distance to the Planck 2018 curve is smaller for the 8-parameter model, an important feature shared by the statistical analysis below. Its inclusion clearly enhances the model’s ability to align more closely with the data, reinforcing the critical role of $\Omega_{\Lambda_{2}}$ in achieving the optimal curve.
@@ 3.2 Analysis by statistical and machine learning techniques
Involving random sampling, the specifics of the brute-force scan may vary slightly with each iteration. It would be reassuring to access the qualitative aspects of the results using an independent method. For this, we apply statistical and machine learning techniques to evaluate the predictive performance of the two cosmological models in estimating the distance variable. The minimum distance predicted by each model is explicitly checked by running CLASS with the corresponding set of the cosmological parameters; see eq. (33) below. Here are some details of the analysis.
@@
The performance of the polynomial regression models is evaluated using $R^{2}$, which quantifies the proportion of variance in the dependent variable modeled by the independent variables. Additionally, the MSE used to measure the average squared difference between predicted and actual values, providing a complementary metric to evaluate model performance. Both models achieve exceptional accuracy, with the $R^{2}$ values nearing 99.9%. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces additional complexity, which increases the data requirements for achieving stabilization on the learning curve. However, this added complexity ultimately enhances the model’s predictive accuracy. The table in Fig. 4 presents a detailed statistical summary of the key variables in the 7-parameter and 8-parameter models. Notably, the addition of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces subtle shifts in other variables, such as $h,n_{s},$ and $\tau_{reio}$. These shifts are accompanied by slightly increased variability in the distance value, as reflected in its higher standard deviation and wider range in the 8-parameter model compared to the 7-parameter model. A series of regression models, ranging from linear regression to polynomial regression up to degree 10, are evaluated for both datasets. The models are built and assessed separately for the 7-parameter and 8-parameter datasets. The quartic regression model (polynomial degree = 4) emerged as the best-performing one based on $R^{2}$ and MSE, as it consistently explains the largest proportion of variance in both the training and testing datasets, compared with higher-degree polynomials that exhibit overfitting tendencies.
Figure 4: Statistical summary of variables in the 7-parameter and 8-parameter models
@@
The $R^{2}$ metric is calculated separately for the training and testing datasets to evaluate the models’ ability to generalize. For the 7-parameter model, the quartic regression achieves an $R^{2}$ of 99.90% on the training set and 99.88% on the testing set, with MSE values of 0.0713 and 0.0795, respectively. In the 8-variable model, including $\Omega_{\Lambda_{2}}$ improved the training $R^{2}$ to 99.92%, though the testing $R^{2}$ remained at 99.88%, with MSE values of 0.0792 and 0.1096. While the addition of $\Omega_{\Lambda_{2}}$ slightly increases the error on the testing set, overall performance remained strong. Figs. 5 and 6 visually confirm these results, showing the predicted versus actual values and residual distributions for both models. The plots reveal a near-perfect alignment along the diagonal.
Figure 7: Optimized feature values for minimum distance in 7- and 8- parameter models
@@
The learning curves for the 7-parameter and 8-parameter models, shown in Figs. 8 and 9, provide valuable insights into how training and validation accuracy evolve as the number of training samples increases. Both models ultimately achieve high accuracy, with training and validation accuracies converging near 99%, highlighting the effectiveness of the quadratic regression approach in predicting the distance variable.For the 7-parameter model, stability is reached around 1,000 training samples, where both accuracies consistently approach 99%. In contrast, the 8-parameter model, which includes the additional parameter $\Omega_{\Lambda_{2}}$, requires a significantly larger dataset to stabilize. Stability is achieved at approximately 2,500 training samples, with both accuracies converging near 99%. This extended stabilization period reflects the added complexity introduced by the extra parameter, which necessitates a larger dataset for effective generalization. Despite the inclusion of $\Omega_{\Lambda_{2}}$, key parameters such as $A_{s}$ and $\omega_{b}$ remain highly stable across both models, demonstrating their robustness to changes in model complexity. The wider range and increased variability in the 8-parameter model underscore its ability to capture more aspects of cosmological phenomena while maintaining the reliability of the core parameters.
@@
The regression results show that the 7-parameter and 8-parameter models predict minimum distance values of 27.88 and 24.39 (rounded), respectively; see the table in Fig. 7. While the 8-parameter model achieves a lower predicted minimum distance, the inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model led to adjustments in other feature values, most notably $\Omega_{\Lambda_{3}}$, which influenced the predicted distance outcome. To validate these predictions, we calculated the distances for the corresponding parameters values; the actual CLASS-computed distance values turn out to be
@@
where the values have been rounded. This result demonstrates a close alignment with the polynomial regression model’s predictions, and confirms that the 8-parameter model achieves a slightly lower minimum distance.

### v2

@@ 3 Modified CLASS and machine learning techniques
In this section, we detail our approach to parameter inference, which fundamentally diverges from the standard Bayesian methods employed by collaborations like Planck (2018). Our methodology emphasizes model building over rigorous parameter inference and focuses on constructing a phenomenological model with additional parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ that can accurately reproduce the observed cosmic microwave background (CMB) angular power spectrum. We perform two complementary analyses: a brute-force parameter scan and a subsequent analysis leveraging statistical and machine-learning techniques. A core component of our approach is the use of a finite-temperature modified version of the cosmological code CLASS. This allows us to perform almost the entire parameter exploration within its Python wrapper, CLASSy, significantly enhancing computational efficiency.
The primary motivation for our non-standard approach is the behavior of the newly introduced parameters, specifically $\Omega_{\Lambda_{2}}$. Initial theoretical estimates for its magnitude were found to be several orders of magnitude larger than the values that ultimately provided the best fit to the Planck 2018 data. Using these initial estimates resulted in power spectra that deviated significantly from the observed data. Instead of a standard Bayesian inference, which would be computationally prohibitive due to the vast parameter space and the lack of a reliable prior for $\Omega_{\Lambda_{2}}$, we adopted an iterative, trial-and-error procedure. This process involved gradually reducing the value of $\Omega_{\Lambda_{2}}$ and performing extensive parameter scans until a good fit was achieved. Although this method lacks statistical rigor (see the conclusion for further discussion), it proved to be an effective strategy for finding the correct order of magnitude for this parameter.
For our analysis, we used the unbinned theoretical comparison spectrum file from the Planck 2018 collaboration:
COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt.
[…]
A key advantage of our methodology is its computational efficiency. The entire parameter search process is performed within CLASSy, with only a small, nearly instantaneous external code required to calculate and sort distance values. In contrast, a standard Bayesian analysis would necessitate additional software packages like MontePython or Cobaya. Our tests showed that integrating these packages can increase runtime by a factor of six to eight. Given that our data collection alone took over three months, this efficiency was critical. While a standard Bayesian approach might require a supercomputer for a similar task, our method provides a more accessible and rapid way to explore this complex parameter space. Once our trial-and-error method establishes a rough order of magnitude for $\Omega_{\Lambda_{2}}$, a more focused scan using tools like Cobaya becomes a feasible and less time-consuming next step.
@@
$H=7.204\times 10^{-19}\,T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}\,T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}+\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}+\frac{2.725}{T}\Omega_{\Lambda_{3}}h^{2}}.$
Once the values of the cosmological parameters are fixed, the power spectra are computed using the modified CLASS. A large number of parameter sets can be scanned by running the CLASS Python wrapper, CLASSy. Each resulting spectrum is then fed into a code that calculates the Euclidean distance between the generated curve and the Planck data [3], COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt. By these comprehensive scans and distance calculations we obtain tens of millions of data points. The details of the brute-force scan will vary slightly each time because the data collection procedure involves random sampling. It will be nice to check the qualitative features of the results by an independent method. For this purpose we employ machine learning techniques and evaluate the predictive performance of the models in estimating the distance.
[…]
The objective is to determine which model best explains the variance in the target variable, the distance $d$, while balancing model complexity and performance. Model selection is based on maximizing $R^{2}$ and minimizing the MSE in both training and testing datasets. Using quartic regression - which proved to be the best-performing among the models we tested, [^10] Regression likely performs well because the distance can, in principle, be expressed as an analytic function of the cosmological parameters. The regression model effectively approximates its multivariable Taylor expansion. including random forest and others - we compare the accuracy and generalization capability of both approaches. The comparison shows that including the additional parameters from finite-temperature QG effects modifies the values of the cosmological parameters.
Variable Count 7-Variable Model ($\Omega_{K}$, 6 variables) 8-Variable Model ($\Omega_{\Lambda_{2}}$, $\Omega_{\Lambda_{3}}$, 6 variables)
Mean STD Range Mean STD Range
distance 5000 55.75 8.19 [28.12, 66.92] 58.80 9.70 [26.83, 71.76]
$\Omega_{\Lambda_{2}}$ 5000 -2.07E-08 2.32E-11 [-2.075E-08, 2.068E-08]
$\Omega_{\Lambda_{3}}$ 5000 0.00233 0.000003 [0.00233, 0.002342] 0.004289 0.00002 [0.00425, 0.00431]
$h$ 5000 0.676876 0.000624 [0.676, 0.678] 0.675491 0.000325 [0.675, 0.676]
$\omega_{b}$ 5000 0.022376 0.000027 [0.02232, 0.02242] 0.022393 0.000028 [0.02235, 0.02245]
$\omega_{cdm}$ 5000 0.119924 0.000207 [0.11944, 0.12033] 0.119853 0.00018 [0.11944, 0.12033]
$A_{s}$ 5000 2.10E-09 3.58E-12 [2.09E-09, 2.1056E-09] 2.10E-09 2.98E-12 [2.0967E-09, 2.1100E-09]
$n_{s}$ 5000 0.965761 0.000491 [0.965, 0.967] 0.965896 0.000487 [0.965, 0.967]
$\tau_{reio}$ 5000 0.053948 0.000757 [0.052, 0.055] 0.054819 0.000591 [0.054, 0.056]
Table 1: Statistical summary of variables in the 7-parameter and 8-parameter models
@@ 3.1 Brute-force scan by CLASSy
In the analysis by employing CLASS, we first conduct brute-force scans to systematically explore the parameter space for cosmological models. By generating discrete power spectra over the multipole moment parameter $\ell$ for suitable ranges of the parameters, their distances from the Planck 2018 experimental curve can be calculated. This approach requires extensive scanning of the parameter space, ultimately determining the set of cosmological parameters that minimizes the distance to the Planck 2018 curve.
As stated in the beginning, the adoption of a distance-based parameter fit is partly motivated by the challenge of determining an appropriate range for $\Omega_{\Lambda_{2}}$. Given specific numerical values for $\Omega_{\Lambda_{2}}$ (along with other parameters), one can efficiently and deterministically compute the distance to the Planck power spectrum curve. This contrasts with the Planck analysis framework - such as that implemented via MontePython - which typically relies on stochastic sampling methods (e.g., MCMC), requiring significantly more computational effort for parameter space exploration. Furthermore, to enhance the robustness and efficiency of our analysis, we incorporate machine learning techniques to guide and refine the parameter search, enabling a more targeted and accelerated convergence toward the best-fit regions.
Using CLASS, specifically its Python wrapper CLASSy, we conduct extensive parameter scans, ultimately collecting several tens of millions of datasets. This immense volume of data is necessary to perform various checks and validate the robustness of the results. Based on these preliminary explorations, appropriate parameter ranges are carefully refined. From these final intervals, we collect 400,000 datasets for the 7-parameter case and 300,000 datasets for the 8-parameter case. The number of the datasets for the 8-parameter model was deliberately chosen less than that of the 7-parameter model. This was to bring out the fact that the former performs better in spite of this disadvantage. See the related comments below.
The impact of inclusion of $\Omega_{\Lambda_{2}}$ is reflected on the actual minimum values of the distance and the corresponding parameters set. For the 7-parameter model, the minimum distance of the 400,000 datasets is [^11] To provide a sense of the accuracy associated with a distance value around 30, consider the following example. Using the CLASS configuration file base_2018_plikHM_TTTEEE_lowl_lowE_lensing.ini, and changing the maximum multipole $\ell$ from 2500 to 2508, yields a computed distance of 56.1 (rounded) when compared to the reference data in COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt. Since base_2018_plikHM_TTTEEE_lowl_lowE_lensing.ini is designed to reproduce the Planck 2018 baseline results, the CLASS-generated spectrum should closely match the Planck theoretical spectrum. The observed discrepancy may arise from factors such as numerical precision, CLASS version, or calibration differences.
@@
$\Omega_{\Lambda_{2}}$ $\Omega_{\Lambda_{3}}$ $h$ $\omega_{b}$ $\omega_{\text{cdm}}$ $A_{s}$ $n_{s}$ $\tau_{\text{reio}}$ $2.073444\times 10^{-8}$ 0.004297 0.675500 0.022394 0.119889 $2.101111\times 10^{-9}$ 0.966000 0.054444
These minima are actual data values. Note the negative sign of $\Omega_{\Lambda_{2}}$ to which we will come back later. Below we obtain the regression functions of these two datasets and obtain the minima predicted by the functions. Those minima will in turn be checked against the values (given in (33)) yielded by directly running CLASS with the corresponding values of the parameters as input. To provide some context for the distance values, we show two plots in Fig. 3. As seen in these plots, the curves with distances in the range of several hundreds are nearly identical to the Planck 2018 curve, appearing directly on top of it.
(a)                    (b)
@@
Given that the 8-parameter model includes one additional parameter, a fair comparison would require a dataset larger than 400,000 by some appropriate factor to adequately cover the expanded parameter space. Despite this discrepancy, the results indicate that the minimum distance to the Planck 2018 curve is smaller for the 8-parameter model, an important feature shared by the statistical analysis below.
@@ 3.2 Analysis by statistical and machine learning techniques
Involving random sampling, the specifics of the brute-force scan may vary slightly with each iteration. It would be reassuring to access the qualitative aspects of the results using an independent method. For this, we apply statistical and machine learning techniques to evaluate the predictive performance of the two cosmological models. The minimum distance predicted by each model is explicitly checked by running CLASS with the corresponding set of the cosmological parameters; see eq. (33) below. Here are some details of the analysis.
@@
The performance of the polynomial regression models is evaluated using $R^{2}$, which quantifies the proportion of variance in the dependent variable modeled by the independent variables. Additionally, the MSE used to measure the average squared difference between predicted and actual values, providing a complementary metric to evaluate model performance. Both models achieve exceptional accuracy, with the $R^{2}$ values nearing 99.9%. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces additional complexity, which increases the data requirements for achieving stabilization on the learning curve. However, this added complexity ultimately enhances the model’s predictive accuracy. Table 1 presents a detailed statistical summary of the key variables in the 7-parameter and 8-parameter models. Notably, the addition of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces subtle shifts in other variables, such as $h,n_{s},$ and $\tau_{reio}$. These shifts are accompanied by slightly increased variability in the distance value, as reflected in its higher standard deviation and wider range in the 8-parameter model compared to the 7-parameter model. A series of regression models, ranging from linear regression to polynomial regression up to degree 10, are evaluated for both datasets. The models are built and assessed separately for the 7-parameter and 8-parameter datasets. The quartic regression model (polynomial degree = 4) emerged as the best-performing one based on $R^{2}$ and MSE, as it consistently explains the largest proportion of variance in both the training and testing datasets, compared with higher-degree polynomials that exhibit overfitting tendencies.
To further assess the robustness and reliability of our regression-based parameter estimates, we apply a non-parametric statistical technique - bootstrap resampling - as a complementary approach to standard inference methods. While our methodology is primarily a model-building exercise rather than a rigorous parameter inference framework, bootstrap analysis offers an empirical means to quantify uncertainty in the optimized parameters and predicted outcomes. From the resulting distributions, we compute empirical 95% confidence intervals (CI) using the 2.5th and 97.5th percentiles, providing a data-driven estimate of sampling variability. These CIs, though not Bayesian credible intervals, serve as practical indicators of the stability and precision of our regression outputs.
@@
3.2.1 Single run results
The $R^{2}$ metric is calculated separately for the training and testing datasets to evaluate the models’ ability to generalize. For the 7-parameter model, the quartic regression achieves an $R^{2}$ of 99.90% on the training set and 99.88% on the testing set, with MSE values of 0.0713 and 0.0795, respectively. In the 8-variable model, including $\Omega_{\Lambda_{2}}$ improved the training $R^{2}$ to 99.92%, though the testing $R^{2}$ remained at 99.88%, with MSE values of 0.0792 and 0.1096. While the addition of $\Omega_{\Lambda_{2}}$ slightly increases the error on the testing set, overall performance remained strong. Figs. 4 and 5 visually confirm these results, showing the predicted versus actual values and residual distributions for both models. The plots reveal a near-perfect alignment along the diagonal.
@@ 3.2.1 Single run results
The learning curves for the 7-parameter and 8-parameter models, shown in Figs. 6 and 7, provide valuable insights into how training and validation accuracy evolve as the number of training samples increases. Both models ultimately achieve high accuracy, with training and validation accuracies converging near 99%, highlighting the effectiveness of the quadratic regression approach in predicting the distance variable. For the 7-parameter model, stability is reached around 1,000 training samples, where both accuracies consistently approach 99%. In contrast, the 8-parameter model, which includes the additional parameter $\Omega_{\Lambda_{2}}$, requires a significantly larger dataset to stabilize. Stability is achieved at approximately 2,500 training samples, with both accuracies converging near 99%. This extended stabilization period reflects the added complexity introduced by the extra parameter, which necessitates a larger dataset for effective generalization. Despite the inclusion of $\Omega_{\Lambda_{2}}$, key parameters such as $A_{s}$ and $\omega_{b}$ remain highly stable across both models, demonstrating their robustness to changes in model complexity.
@@
The regression results show that the 7-parameter and 8-parameter models predict minimum distance values of 27.88 and 24.39 (both rounded), respectively; see Table 2. We again note that the sign of $\Omega_{\Lambda_{2}}$ is negative; the same sign is observed in [27], where its physical significance is discussed. While the 8-parameter model achieves a lower predicted minimum distance, the inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model led to adjustments in other feature values, most notably $\Omega_{\Lambda_{3}}$, which influenced the predicted distance outcome. To validate these predictions, we calculated the distances for the corresponding parameters values; the actual CLASS-computed distance values turn out to be
distance of 7-parameter model distance of 8-parameter model
26.03 24.95
@@
where the values have been rounded. This result confirms that the 8-parameter model achieves a slightly lower minimum distance.
Variable 7-Variable Model 8-Variable Model
$\Omega_{\Lambda}$ - 2.077437E-08
$\Omega_{\Lambda}/\Omega_{\Lambda 3}$ 0.00233471 0.00432285
$h$ 0.67733813 0.67537001
$\omega_{b}$ 0.02238564 0.02238631
$\omega_{cdm}$ 0.11984093 0.11985634
$A_{s}$ 2.094146E-09 2.09484E-09
$n_{s}$ 0.96599393 0.96583363
$\tau_{reio}$ 0.05295666 0.0530449
predicted distance 27.877821 24.391195
Table 2: Optimized feature values for minimum distance in 7- and 8- parameter models
3.2.2 The results with bootstrap resampling
Variable 7-Variable Model (Mean [95% CI]) 8-Variable Model (Mean [95% CI])
$\Omega_{\Lambda_{2}}$ -2.070386E-08 [-2.073444E-08, -2.068444E-08]
$\Omega_{K}/\Omega_{\Lambda_{3}}$ 2.336606E-03 [2.331111E-03, 2.338889E-03] 4.282391E-03 [4.2700E-03, 4.304444E-03]
$h$ 0.6773487 [0.6773333, 0.6780] 0.6753918 [0.6753333, 0.6755]
$\omega_{b}$ 0.02239074 [0.02238667, 0.02239778] 0.02239451 [0.02239444, 0.02239444]
$\omega_{cdm}$ 0.1198824 [0.1196667, 0.1198889] 0.1198889 [0.1198889, 0.1198889]
$\Lambda_{s}$ 2.098202E-09 [2.096667E-09, 2.101111E-09] 2.099984E-09 [2.098889E-09, 2.103333E-09]
$n_{s}$ $0.9660145$ [$0.9660$, $0.9665$] $0.9659412$ [$0.9658889$, $0.9660$]
$\tau_{reid}$ $0.05385517$ [$0.05350$, $0.0548333$] $0.05422211$ [$0.05400$, $0.05500$]
predicted distance $33.30075$ [$33.08834$, $33.6667$] $33.35167$ [$33.06932$, $33.69121$]
Table 3: Optimized Feature Values and 95% Confidence Intervals
To rigorously assess the variability and robustness of the regression-based optimization process, we incorporate a bootstrap resampling procedure - a widely used non-parametric technique in statistical analysis. This approach allows us to empirically quantify the uncertainty in our best-fit cosmological parameters and predicted distance values, independent of any prior distributional assumptions.
In this analysis, we perform 1,000 bootstrap resampling iterations for both the 7-parameter and 8-parameter datasets. In each iteration, the training set (comprising 70% of the full dataset, i.e., 3,500 samples) is resampled with replacement to generate a new training subset of the same size. The quartic polynomial regression model—identified earlier as the best-performing model based on $R^{2}$ and mean squared error (MSE)—is then refitted using the resampled data. For each refitted model, we identify the parameter configuration that minimizes the predicted distance, effectively simulating the optimization process under perturbed training data conditions. This repeated process yields a distribution of best-fit parameter values and predicted distance outcomes across the 1,000 bootstrap samples.
From these distributions, we compute empirical 95% confidence intervals (CIs) by extracting the 2.5th and 97.5th percentiles of each parameter’s distribution. These intervals provide a direct, data-driven estimate of the sampling variability inherent in our modeling framework. Unlike Bayesian credible intervals, which require specification of prior distributions, bootstrap-derived CIs rely solely on the observed data and model fit, making them especially suitable for complex, high-dimensional systems where formal likelihood functions may be intractable or computationally expensive.
The results of this analysis, summarized in Table 3, reveal consistently narrow confidence intervals for most cosmological parameters in both the 7- and 8-parameter models. This suggests a high degree of robustness in the regression-based optimization process: the estimated values are relatively stable across random perturbations of the input data. The predicted distance values likewise exhibit limited variability under bootstrap resampling.
Interestingly, the inclusion of the additional parameter $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces modest shifts in the distributions of several other parameters - most notably the Hubble parameter $h$, the scalar spectral index $n_{s}$, and the reionization optical depth $\tau_{reio}$. These shifts are accompanied by slightly broader confidence intervals and increased standard deviation in the predicted distance values, indicating a mild increase in model complexity and sensitivity to input variability. Nevertheless, the confidence intervals for these parameters in both models overlap substantially, suggesting that the two regression models produce statistically consistent outputs.
Taken together, the bootstrap results lend strong support to the generalizability and robustness of our quartic regression models. They confirm that the optimized parameter values obtained via our data-driven procedure are not artifacts of a particular sample realization, but instead reflect stable features of the underlying parameter-distance relationship as encoded in the CLASSy-generated datasets.